OpenStudy (anonymous):

567 is what percent of 420

OpenStudy (anonymous):

el panorama es un poco confuso o.o

OpenStudy (anonymous):

567 / 420 = 1.35 1.35*100% = 135% "567 is 135% of 420."

OpenStudy (anonymous):

567 = X/100 * 420 567/420 = X/100 1.35 * 100 = X X = 135%

OpenStudy (anonymous):

U.S.A

OpenStudy (anonymous):

wut ?

OpenStudy (anonymous):

@sebasss if the people helping her can sort of understand what she's saying, then why does it matter to you?

OpenStudy (anonymous):

i never said it matterd lol all i said was u.s.a cause he said it , and yall be trippn

OpenStudy (anonymous):

its alright lol

OpenStudy (anonymous):

people be trippn hard now adays